What is the pH at the equivalence point in the titration of a 28.7 ml. sample of a 0.314 M aqueous hydrofluoric acid solution with a 0.342 M aqueous barium hydroxide solution


Ka=7.2*10^-4

What is the pH at the equivalence point in the titration of a 28.7 ml. sample of a 0.314 M aqueous hydrofluoric acid solution with a 0.342 M aqueous barium hydroxide solution?
Answer #1

HF + Ba(OH)2----> BaF2 + H2O

0.314M. 0.342 M

0. 0.28 M 0.314 M

Finally we get mixture of two bases, so we will calculate pH from strong base Ba(OH)2 here.

Conc. of OH- = 2×0.28=0.56

pOH = -log 0.56 = 0.252

pH= 14-0.252= 13.748
Answer #2

Ba(OH)2 + 2HF Bar + 2tt20 Imole Ba(OH) 2 can neutralize 2 moleste I Mive = 2 M2 V2 > 20% 0.314 x 29.7 - 2x0.342 X V2 0.314 x